The average time for a courier to travel from Tagum to Davao is 120 minutes, and the standard deviation is 15 minutes. If there are 1255 courier, how many takes between 100 to 140 minutes to travel from Tagum to Davao?


1
Expert's answer
2022-03-19T02:35:18-0400

"\mu=120,\sigma=15,n=1255"

"P(100<X<140)=P(\frac{100-120}{15}<Z<\frac{140-120}{15})"

"=P(-1.33<Z<1.33)=PZ<1.33)-P(Z<-1.33)"

"=0.9082-0.0918=0.8164"

"=0.8164*1255=1024.6=1025" couriers